Two forces act on the same object: F1 = 51.2 N at 52.1o and F2 = 80.1 N at 132o. Find the magnitude of the resultant force, in N.
1
Expert's answer
2020-02-05T12:27:06-0500
"F=\sqrt{F_1^2+F_2^2+2F_1F_2\cos{(132-52.1)}}"

"F=\sqrt{51.2^2+80.1^2+2(51.2)(80.1)\cos{(79.9\degree)}}=102\ N"